OpenStudy (anonymous):

Carla drove 180 miles from her home to Huston in 3 hours. On her return home the traffic was slow and the trip took 4 hours. Find Carla's average rate of speed for her drive to Huston,for her return trip,and for the round trip to the nearest mile per hour.

OpenStudy (anonymous):

explanation please :DD

OpenStudy (anonymous):

The average speed is v = d/t where v is speed, d is distance, and t is time. In her drive to Houston, the distance is 180 miles and her time was 3 hours. Thus, her average speed is 180/3 = 60 miles per hour. In her drive from houston to her home, the distance is the same but the time is 4 hours. by the same formula, average speed is 180/4 = 45 miles per hour. The roundtrip is NOT the sum of the two average speeds. We must apply the formula one more time. The total distance is 360. The total time is 7 hours. Thus her average velocity is 360/7 miles per hour.

OpenStudy (anonymous):

thank you you are very helpful :)